### v2

# Mapple: A Domain-Specific Language for Mapping Distributed Programs

Abstract.
[…]
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Optimizing parallel programs on modern distributed systems remains complex and labor-intensive despite decades of research. Programmers must navigate intricate low-level programming interfaces and often make substantial code modifications to achieve high performance, making tuning challenging even for experts.
Two dominant programming paradigms exist today. The first is the Message Passing Interface (MPI)+X model (gropp1999using; sanders2010cuda; chandra2001parallel; dagum1998openmp; yang2011hybrid) where MPI is used for distributed-memory parallelism (across computers), and X is typically a shared-memory programming model for intra-node parallelism (within one computer), such as OpenMP or CUDA. This model gives programmers fine-grained control over performance-critical aspects but significantly compromises productivity. Achieving performance gains often requires substantial changes throughout the application code, making it difficult to optimize without inadvertently affecting application logic. The second paradigm is task-based programming systems (chamberlain2007parallelchapel; augonnet2009starpu; fatahalian2006sequoia; kale1993charm++; duran2011ompss; kaiser2014hpx; slaughter2015regent), which separate performance decisions from core application logic. Tasks are functions whose dependencies are analyzed by the runtime and scheduled across processors. These systems offer interfaces for specifying performance policies such as task and data placement, allowing developers to tune performance independently of application logic. This separation improves modularity and makes tuning more systematic and less error-prone.
[…]
A major limitation of existing mapping interface designs is their low-level nature and lack of abstraction, which exposes users to the complexity and intricacies of the underlying runtime system. These interfaces are often tightly coupled with the runtime, making them difficult to use, especially for application developers unfamiliar with runtime internals. For instance, defining a new mapper in the Chapel framework (chamberlain2007parallelchapel) requires implementing 19 functions according to 33 pages of documentation (chamberlain2010user; chamberlain2011authoring); the Legion mapping interface has more user-implementable callbacks (bauer2012legion).
[…]
A core challenge in any mapping interface targeting distributed applications is to determine how a multi-dimensional task space, defined by the algorithm’s set of tasks to execute, maps onto a multi-dimensional processor space, defined by the machine’s hierarchical architecture (e.g., a cluster of $n$ nodes with $m$ GPUs per node). This mapping process is directly responsible for the amount of data movement across processors, which is a key factor in achieving high performance on distributed systems. This challenge is compounded by the fact that these two spaces often differ in dimensionality, making writing an effective mapping inherently complex.
[…]
While inspired by array programming systems such as APL (iverson1962programming), ZPL (lin1993zpl), and NumPy (harris2020array), Mapple extends these ideas to distributed task-based programs. The processor space transformation is central to the decompose primitive, which has no counterpart in prior work. Moreover, Mapple provides a richer set of mapping decisions, including processor selection, memory placement, data layout, garbage collection, and load balancing (see Section 7). These capabilities enable Mapple to support a broader range of distributed applications, including tensor algebra and scientific simulations, going beyond prior distributed scheduling systems such as Distributed Halide (denniston2016distributed), which are restricted to stencil computations and CPU-only environments.
While Mapple’s transformation primitives resemble classical loop transformations such as tiling, fusion, and fission (wolf1992improving; chen2008framework; ragan2013halide; chen2018tvm), they differ in purpose and scope. Loop transformations schedule loops, whereas Mapple schedules sets of tasks across distributed machines. Likewise, Mapple shares the goal of separating performance decisions from algorithms with prior scheduling languages (zhang2018graphit; senanayake2020sparse; ahrens2022autoscheduling), but these do not handle dimensionality mismatches between index spaces or explicitly minimize inter-processor data movement.
[…]
In summary, the contributions of this work are:
  1. (1)
    A high-level programming interface for mapping distributed applications.
  2. (2)
    A set of transformation primitives for resolving dimensionality mismatches between task and processor spaces, including a key primitive, decompose, which helps minimize data communication volume.
  3. (3)
[…]
  4. (4)
[…]
2. Design Goals and Non-Goals
m = Machine(GPU)
def block2d(Tuple point, Tuple space):
idx = point * m.size / space
return m[*idx]
IndexTaskMap tiles block2d
Region task_init * GPU FBMEM
Layout task_finish * CPU C_order
GarbageCollect systolic *
Backpressure systolic 1
(a) A Mapple mapper.
ShardID shard(const DomainPoint& point,
const Domain& space,
const size_t num_nodes) {
auto rect2d = Rect<2>(space);
// Implementation of external helper function is omitted
auto blocks = this->select_num_blocks<2>(num_nodes, rect2d);
Point<2> zeroes{0}, ones{1};
Rect<2> blockSpace(zeroes, blocks - ones);
auto numPoints = rect2d.hi - rect2d.lo + ones;
Point<2> projected = point * blocks / numPoints;
AffineLinearizedIndexSpace<2> linearizer(blockSpace);
return linearizer.linearize(projected);
void slice(const Task& task,
const SliceTaskInput &input,
SliceTaskOutput &output) {
vector<Processor> targets = this->select_targets(task);
auto size = targets.size();
DomainT<2> space = input.domain;
Point<2> n_b = this->select_num_blocks<2>(size, space);
Point<2> zeros{0}, ones{1};
... // 125 lines of C++ code omitted here
for (PointInRectIterator<2> it(blocks); it()!=NULL; it++) {
Point<2> lo = *it; Point<2> hi = *it + ones;
Point<2> slice_lo = n_p * lo / n_b + space.lo;
Point<2> slice_hi = n_p * hi / n_b + space.lo - ones;
DomainT<2,coord_t> slice_space; TaskSlice slice;
slice.domain = {slice_lo, slice_hi};
slice.proc = targets[index++ % targets.size()];
output.slices.push_back(slice);
(b) Simplified code excerpts from a C++ mapper.
Figure 1. Comparison between a Mapple mapper and its partial C++ counterpart. The Mapple mapper uses a high-level, declarative design that abstracts away the complexity of low-level C++ implementations while still supporting performance optimization. The boxed block2d function is realized through two separate APIs in the C++ mapper, illustrating the conciseness of Mapple.
Our goal is to provide a high-level programming interface that abstracts away the complexities of the underlying runtime, allowing users to focus on mapping logic rather than system internals. Specifically, Mapple enables users to place tasks across distributed machines, determine where data resides in memory, control data layout, manage garbage collection, and make scheduling decisions. We introduce abstractions and transformation primitives that make such decisions directly programmable and easy to change. For example, block2d specifies task space distribution, Region and Layout control data placement and layout, GarbageCollect handles memory management, and Backpressure guides scheduling. The Mapple mapper in Figure 1 achieves the same functionality as the C++ implementation from prior work (yadav2022distal) but with significantly less code and greater clarity. We quantify the reduction in code size in Section 6.1.
Figure 1(a) presents only a subset of Mapple’s features. We present the complete feature set of Mapple in Section 7. The effectiveness of Mapple for performance tuning is demonstrated in Section 6.2. We provide a full Mapple mapper in the supplementary material (Section A.1).
We do not aim to support portability across different runtime systems. This non-goal reflects the inherent differences among distributed runtimes, which often prevent direct reuse of implementations. Nonetheless, we believe the ideas and abstractions introduced here can inform the design of mapping interfaces for other distributed systems.
3. A Core Challenge: Mapping Task Spaces to Processor Spaces
A key challenge for any mapping interface in distributed applications is assigning tasks to processors. The algorithm defines a multi-dimensional task space consisting of indexed tasks, while the hardware defines a multi-dimensional processor space, capturing its hierarchical structure (e.g., a cluster of $n$ nodes with $m$ GPUs per node). The core problem, index mapping, is to define a function mapping task indices in the task space to processor indices in the processor space that minimizes data movement, which is crucial for high performance on distributed systems. Differences in dimensionality between the two spaces further complicate this task.
Task-based systems vary in how they define task spaces. Some do it implicitly (e.g., through iteration spaces of loops) and some do it explicitly (e.g., Legion’s index tasks). We assume we are given one or more task spaces to map.
We first survey existing approaches to index mapping. Motivated by their limitations, we introduce Mapple’s transformation primitives designed for index mapping in distributed systems. We defer the discussion of how Mapple handles dimensionality mismatches between task and processor spaces, using the decompose primitive, to Section 4.
Figure 2. Three existing interface designs for mapping task space to processor space: the enumeration-based, keyword-based, and programmatic approaches.
3.1. Existing Approaches
Existing task-based systems support mapper development through one of three design styles, as illustrated in Figure 2. The enumeration-based interface (ampi) allows users to explicitly specify how each tile of the distributed array is mapped to a processor. While this interface is expressive, it lacks generality: enumeration-based mappers are hard-coded to a specific size of inputs or machines.
In contrast, keyword-based designs (keyword) provide a higher-level abstraction by letting users choose from a predefined set of standard distributions (e.g., BlockDist for block distribution). However, this approach sacrifices flexibility—the fixed keyword-based options cannot express non-standard mapping strategies. For example, none of the distributions used in the matrix-multiplication algorithms in Section 6.1 can be represented using keyword-based interfaces.
The programmatic approach defines an interface of classes and methods that must be implemented to control mapping behavior. While this design offers flexibility, existing programmatic interfaces expose significant system detail and complexity. For instance, defining a new mapping in Chapel requires implementing 19 functions based on 33 pages of documentation (chamberlain2010user; chamberlain2011authoring), and Legion’s programmatic interface directly exposes the runtime’s internal abstractions, accompanied by 19 pages of documentation (bauer2012legion). In our experience, only users with substantial expertise in the underlying runtime system are able to successfully develop mappers using such interfaces.
3.2. Motivating Examples
Figure 3. Block mapping from the task space $(6,6)$ to the processor space $(2,2)$, a machine with 2 nodes and 2 GPUs per node. A node index and a GPU index within the node name a specific GPU processor. The shaded index point $(2,3)$ is mapped to node 0 and GPU 1.
We present three example mappers in Mapple to illustrate how task spaces are mapped onto processor spaces. The task space comes from the application code, which is separate from the mapper. We show how tasks and task spaces are defined in Section A.1 of the supplementary material.
Below, we describe a standard block distribution in Section 3.2.1 to introduce the core concepts. Next, we discuss a custom cyclic distribution in Section 3.2.2, which is not supported in keyword-based mapping interfaces. Finally, we use a mapper from Solomonik’s 2.5D matrix multiplication algorithm (solomonik2011communication) in Section 3.2.3 to demonstrate the complexity introduced by mismatched dimensionality between the task and processor spaces.
3.2.1. Standard Block Distribution
Block distribution is a common default strategy widely supported by parallel programming systems. We use it here to illustrate the core concepts of index mapping in Mapple. In Mapple, users define mappers by writing a function that maps each point in the task space to a target processor in the processor space. Figure 3 shows the function block2D, which implements a block distribution for a task space of size $(6,6)$ (that is, $0\leq x<6$ and $0\leq y<6$) and a processor space of size $(2,2)$, representing two nodes with two GPUs per node. The function assigns each point to a specific processor.
In Mapple, the GPU processor space can be accessed by Machine(GPU), which is a 2D tuple whose shape, representing the number of nodes and GPUs per node, is automatically inferred by the runtime. Consequently, users do not need to specify these dimensions explicitly, allowing Mapple mappers to remain portable across machines of different sizes. In the mapper function, the index point ipoint, task space ispace, and processor space size m.size are all 2D integer tuples. These values are used to compute the processor indices node_idx and gpu_idx, which specify the destination processor for each index point. As shown in Figure 3, the shaded point $(2,3)$ is mapped to node 0 and GPU 1. The function returns a specific processor for each index point, resulting in a block distribution over the processor space.
Figure 4. A custom cyclic distribution. The merge primitive transforms the 2D processor space into a 1D space. The mapping function linearizes each 2D index point and applies a round-robin distribution over the resulting 1D processor space.
3.2.2. Custom Cyclic Distribution
We present a custom distribution expressible in Mapple but unsupported by most parallel programming systems that limit users to predefined keyword-based schemes. This non-standard cyclic distribution, inspired by a simplified variant of a matrix multiplication algorithm, is implemented by the mapping function linearCyclic in Figure 4, where the subdiagonal index points (shaded) map to the first processor.
The mapper first applies the merge primitive, which fuses dimensions 0 and 1 of the 2D processor space into a single dimension (formally defined in Section 3.3). The resulting processor space m is one-dimensional with size 4. The 2D task index points are then linearized and assigned to processors in a round-robin fashion.
3.2.3. Mapping Under Dimensionality Mismatch
Figure 5. A mapper illustrating the dimensionality mismatch between task and processor spaces in the Solomonik’s algorithm on a 2-node, 4-GPU-per-node machine. The original 2D processor space is transformed to 6D via the split primitive, shown as two 3D spaces.
We give an example illustrating the complexity introduced by a mismatch between the dimensionality of the task space and the processor space. Figure 5 shows a mapper for the Solomonik’s algorithm executed on a 2-node machine, where each node has 4 GPUs. The task space is three-dimensional, and the algorithm specifies a hierarchical distribution: the task space is first partitioned along the x-axis so that each node handles half of the x-dimension, and within each node, the 4 GPUs perform a 2D block distribution over the y-z plane.
The task space is three-dimensional, while the processor space is initially two-dimensional. To support the mapping required by the algorithm, we apply the split primitive four times (marked in red in the code). The first two splits align the node dimension with the task space’s x-axis, and the last two align the GPU dimensions with the y- and z-axes. This produces a six-dimensional processor space, visualized as two 3D spaces: one for nodes $(2,1,1)$ and one for GPUs $(1,2,2)$. The split primitive is formally defined in Section 3.3.
This example also raises deeper research questions. Can we automatically determine the appropriate splitting factors? What underlying principles guide the mapping decisions when resolving mismatched dimensions? Is it possible to address dimensionality mismatches through a generalized transformation? We explore these questions in Section 4, where we introduce a new primitive called decompose to provide a unified solution.
3.3. Transformation Primitives
Transformation Semantics
$m^{\prime}=m.\mathtt{split}(i,d)$ $m^{\prime}[a_{0},\ldots,a_{n}]\coloneq m[b_{0},\ldots,b_{n-1}]$
$b_{t}=\begin{cases}a_{t}\qquad&\quad t<i\\ a_{i}+a_{i+1}\cdot d\qquad&\quad t=i\\ a_{t+1}\qquad&\quad t>i\end{cases}$
$m^{\prime}=m.\mathtt{merge}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-2}]\coloneqq m[b_{0},\ldots,b_{n-1}]$
[…]
$m^{\prime}=m.\mathtt{swap}(p,q)$ $m^{\prime}[a_{0},\ldots,a_{n-1}]\coloneq m[b_{0},\ldots,b_{n-1}]$
[…]
$m^{\prime}=m.\mathtt{decompose}(i,T)$ $T=(l_{1},\ldots,l_{k})$
Formally defined in Section 4
Figure 6. Semantics of transformation primitives expressed as mappings from the indices of the transformed processor space to the indices of the original processor space.
We define the semantics of each Mapple transformation primitive in Figure 6, with the exception of the decompose primitive, which is deferred to Section 4. Each transformation primitive is a function that takes a processor space $m$ as input and returns a transformed processor space $m^{\prime}$, following the mapping illustrated on the right-hand side of Figure 6. We now describe each transformation.
The split transformation takes two arguments: a dimension index $i$ and a splitting factor $d$. Given a processor space $m$ of shape $(s_{0},\dots,s_{n-1})$, the operation $m^{\prime}=m.\mathtt{split}(i,d)$ produces a new processor space $m^{\prime}$ of shape $(s_{0},\dots,d,s_{i}/d,\dots,s_{n-1})$. Like other processor space transformations, this transformation is invertible; in our implementation, Mapple uses the inverse to translate from mapper operations on $m^{\prime}$ to the original space $m$.
The merge transformation takes two dimensions $p$ and $q$ of the processor space and fuses them into one. Given a processor space $m$ of shape $(s_{0},\dots,s_{n-1})$, the operation $m^{\prime}=m.\mathtt{merge}(p,q)$ produces a new processor space $m^{\prime}$ of shape $(s_{0},\dots,s_{p}\cdot s_{q},\dots,s_{n-1})$, where the two dimensions at positions $p$ and $q$ in $m$ are combined into a single dimension at position $p$ in $m^{\prime}$. The index mapping is given by: $m^{\prime}[a_{0},\dots,a_{n-2}]=m[a_{0},\dots,a_{p}\bmod s_{p},\dots,\lfloor a_{p}/s_{p}\rfloor,\dots,a_{n-2}]$.
Transformation primitives in Mapple can be composed sequentially. Consider a 2D processor space $m$, and let $m^{\prime}=m.\mathtt{split}(0,d)$, followed by $m^{\prime\prime}=m^{\prime}.\mathtt{merge}(0,1)$. The resulting processor space $m^{\prime\prime}$ is again 2D. We now derive the index transformation from $m^{\prime\prime}$ to the original space $m$ by sequentially applying the semantics of the merge and split transformations: $m^{\prime\prime}[a_{0},a_{1}]=m^{\prime}[a_{0}\mathbin{\bmod}d,\lfloor a_{0}/d\rfloor,a_{1}]=m[(a_{0}\mathbin{\bmod}d)+\lfloor a_{0}/d\rfloor\times d,a_{1}]$. Because the expression $(a_{0}\mathbin{\bmod}d)+\lfloor a_{0}/d\rfloor\times d$ simplifies to $a_{0}$ for all integers, it follows that $m^{\prime\prime}[a_{0},a_{1}]=m[a_{0},a_{1}]$merge is the inverse of split.
The swap transformation exchanges two dimensions of the processor space. Combined with merge, which flattens two dimensions into one, users can control whether merging follows row-major or column-major order.
Using these primitives, Mapple can transform processor spaces in flexible ways. Figure 7 illustrates how Mapple supports common distribution patterns. Suppose we want to map a 2D task space onto a 2-node machine with 2 GPUs per node. In the figure, the shaded region of the task space is mapped to the corresponding shaded processor in the processor space.
For the block2D distribution, the mapping function is equivalent to the one in Figure 3, but is simpler thanks to Mapple’s support for tuple arithmetic. Tuple operands must have the same shape. In Figure 7, both the transformed processor spaces and the task space are 2D. In particular, the * operator flattens tuples for indexing: for example, m[*(1,2)] is equivalent to m[1,2]. Mapple supports both elementwise arithmetic on individual tuple dimensions and operations over entire tuples, enabling users to express index mappings more concisely and intuitively.
Figure 7. Common distributions expressed in Mapple. The shaded region in the task space is mapped to the corresponding shaded processor. The transformation code reshapes the original $(2,2)$ processor space into the desired processor space, which is then used in the user-defined mapping function.
4. Decompose Transformation Primitive
As discussed in Section 3.2.3, mismatches between the dimensionality of the task and processor spaces pose significant challenges for mapping decisions. In this section, we address this issue by analyzing its underlying principles and introducing a generalized transformation primitive, decompose. Section 4.1 analyzes a suboptimal, but standard, heuristic to handle dimensionality mismatches. We then formally define the decompose primitive in Section 4.2, grounded in a communication volume analysis. Finally, Section 4.3 describes our search-based optimization algorithm and analyzes its complexity.
4.1. Suboptimal Existing Heuristics for Resolving Dimensionality Mismatch
We examine how existing task-based systems address dimensionality mismatches between task and processor spaces. Most frameworks sidestep this issue by linearizing both spaces and applying a default 1D block mapping. However, they rarely document their linearization procedures, making direct comparison difficult. An exception is Chapel (chamberlain2007parallelchapel), a widely used parallel programming framework, which explicitly describes its approach. To expose the limitations of this strategy, we present an example showing how it can lead to suboptimal mappings.
Suppose we have 6 processors and a 2D task space. To match the dimensionality, we must split the 6 processors into a 2D tuple. There are four possible factorizations: $(6,1)$, $(3,2)$, $(2,3)$, and $(1,6)$. The existing heuristic used to determine the processor grid, as shown in Algorithm 1 of the supplementary material, does not consider the shape of the task space.
The heuristic follows a greedy strategy, aiming to produce factors that are as balanced as possible in magnitude. In the example above, the algorithm selects the grid $(3,2)$ by prioritizing balanced factorization. Now consider two task spaces, $(12,18)$ and $(18,12)$, for an application with nearest neighbor locality (i.e., data is exchanged among immediate neighbors in task space), as shown in Figure 8. Each task space is mapped to the processor grid $(3,2)$ using the block2D function, which partitions the task space into rectangular blocks assigned to each processor. The orange regions in the figure indicate data that must be transferred across processor boundaries due to this mapping.
For the $(12,18)$ task space, the total inter-processor communication volume is 96 elements, while for the $(18,12)$ task space it is 84 elements. This difference highlights the sensitivity of communication cost to how the task space is aligned with the processor grid. If the algorithm had taken the shape of the task space into account and chosen the processor grid $(2,3)$ for the $(12,18)$ space, the communication volume could have matched the more efficient 84-element case, avoiding unnecessary communication overhead.
Figure 8. For this application, the mapper from Algorithm 1 (see supplementary material) selects a fixed $(3,2)$ processor grid for 2D block mapping. Orange regions indicate inter-processor data transfers. The $(12,18)$ task space incurs higher communication than $(18,12)$, showing that the chosen grid is suboptimal; a $(2,3)$ grid would yield lower communication.
4.2. Formal Definition of Decompose and Intuitive Solution
We formally define the optimization problem that the decompose primitive aims to solve, based on an analysis of inter-processor communication volume. We also present the underlying intuition to clarify how the solution addresses this problem. Implicit in all solutions to this problem (including the greedy heuristic described above) are two assumptions: that tasks in the task space have uniform cost (i.e., are load balanced) and that there is locality in the communication pattern, so that partitioning the task space into compact regions is desirable to minimize communication.
Suppose we apply decompose to a processor space $M$ of size $(d_{0},\ldots,d_{n-1})$, written as $M^{\prime}=M.\texttt{decompose}(i,\allowbreak(l_{1},\ldots,l_{k}))$. The decompose primitive splits the $i$-th dimension $d_{i}$ into $k$ natural numbers $d_{i_{1}},\ldots,d_{i_{k}}$, producing a new processor space $M^{\prime}$ of size $(d_{0},\ldots,d_{i_{1}},\allowbreak\ldots,\allowbreak d_{i_{k}},\allowbreak\ldots,d_{n-1})$. The transformation preserves the total size of the processor space, so it must satisfy $\prod_{m=1}^{k}d_{i_{m}}=d_{i}$.
Conceptually, decompose is a shorthand for applying a sequence of split transformations. Specifically, applying $m_{k}=m_{1}.\texttt{decompose}(i,T)$ is equivalent to performing $m_{n+1}=m_{n}.\texttt{split}(i+n-1,d_{i_{n}})$ for $1\leq n<k$. There can be multiple ways to factor $d_{i}$ into $k$ natural numbers $d_{i_{1}},\ldots,d_{i_{k}}$. The decompose primitive selects a factorization by solving the following optimization problem:
$\min_{\{d_{i_{m}}\in\mathbb{N}\}}\sum_{m=1}^{k}\frac{d_{i_{m}}}{l_{m}}\quad\text{s.t.}\quad\prod_{m=1}^{k}d_{i_{m}}=d_{i}.$
An equivalent formulation introduces the workload vector $w_{m}=\frac{l_{m}}{d_{i_{m}}}$ for $1\leq m\leq k$, where $w_{m}$ represents the portion of the task space assigned to each processor in the $m$-th dimension. Rewriting the problem in terms of the workload vector gives:
$\min_{\{w_{m}\}}\sum_{m=1}^{k}\frac{1}{w_{m}}\quad\text{s.t.}\quad\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}},\;\frac{l_{m}}{w_{m}}\in\mathbb{N}.$
Figure 9. We show two examples of how to compute the communication volume (namely, the number of inter-processor elements). In the 2D example, we compute the length of the inter-rectangle line by summing up the perimeter of all 4 rectangles of $(w_{1},w_{2})$ and subtracting the perimeter of the $(l_{1},l_{2})$ rectangle. In the 3D example, we compute the area of the inter-cuboid surface by summing up the surface area of all 16 cuboids of $(w_{1},w_{2},w_{3})$ and subtracting the surface area of the $(l_{1},l_{2},l_{3})$ cuboid.
We now illustrate the optimization problem using the two examples shown in Figure 9. The left-hand side depicts a 2D case with $k=2$ and processor count $d_{i}=4$. To compute the communication volume, we count the number of elements transferred across processor boundaries. This volume equals $2\cdot(w_{1}+w_{2})\cdot d_{i}-2\cdot(l_{1}+l_{2})$, where the first term is the total perimeter of the $d_{i}$ blocks of size $(w_{1},w_{2})$, and the second term is the perimeter of the entire task space of size $(l_{1},l_{2})$.
Given fixed input dimensions $l_{1}$ and $l_{2}$ and processor count $d_{i}$, the second term is constant, and $w_{1}\cdot w_{2}=\frac{l_{1}\cdot l_{2}}{d_{i}}$ is also fixed. Therefore, minimizing the communication volume reduces to minimizing $w_{1}+w_{2}$, which is equivalent to minimizing $\frac{1}{w_{1}}+\frac{1}{w_{2}}$. This matches the objective of our optimization formulation in the 2D case.
On the right-hand side, the figure shows a 3D example with $k=3$ and $d_{i}=16$. Here, the communication volume corresponds to the total inter-block surface area. The surface area $S$ of the $d_{i}$ cuboids of size $(w_{1},w_{2},w_{3})$ is given by $2S=2(w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3})\cdot d_{i}-2(l_{1}l_{2}+l_{1}l_{3}+l_{2}l_{3})$. Given fixed input dimensions $l_{1},l_{2},l_{3}$ and processor count $d_{i}$, the term $w_{1}w_{2}w_{3}=\frac{l_{1}l_{2}l_{3}}{d_{i}}$ is constant. Thus, minimizing $S$ is equivalent to minimizing $w_{1}w_{2}+w_{1}w_{3}+w_{2}w_{3}$. Under the fixed-product constraint, this can be transformed into minimizing $\frac{1}{w_{1}}+\frac{1}{w_{2}}+\frac{1}{w_{3}}$, again aligning with the objective of our optimization formulation in the 3D case. This analysis generalizes to any $k$-dimensional case. Let $S$ denote the total inter-surface area in the $k$-dimensional setting. Then,
$2S=SA(w_{1},w_{2},\ldots,w_{k})\cdot d_{i}-SA(l_{1},l_{2},\ldots,l_{k}),$
where $SA(x_{1},x_{2},\ldots,x_{k})$ denotes the surface area of a $k$-dimensional hyperrectangle, defined as
$SA(x_{1},x_{2},\ldots,x_{k})=2\cdot\left(\prod_{m=1}^{k}x_{m}\right)\cdot\left(\sum_{m=1}^{k}\frac{1}{x_{m}}\right).$
Given that $\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$ is constant for fixed input size and processor count, minimizing $S$ reduces to minimizing $\sum_{m=1}^{k}\frac{1}{w_{m}}$—which matches the objective of the optimization problem introduced earlier.
The solution to the optimization problem is not immediately obvious. In the following, we present an intuitive justification for the solution, which is based on the classic inequality stated below.
Theorem 0.
Given a list of $n$ positive numbers $a_{1},a_{2},\ldots,a_{n}$, the following inequality holds:
$\frac{1}{n}\sum_{m=1}^{n}a_{m}\geq\left(\prod_{m=1}^{n}a_{m}\right)^{1/n},$
with equality if and only if $a_{1}=a_{2}=\cdots=a_{n}$.
This is the well-known arithmetic-geometric mean (AM-GM) inequality. It states that the arithmetic mean of a set of positive numbers is always greater than or equal to their geometric mean, with equality only when all values are equal. In the optimization problem, the term $\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$ is a constant determined by the input size and the number of processors. We can now apply the inequality to the objective: $\frac{1}{k}\sum_{m=1}^{k}\frac{1}{w_{m}}\geq\left(\prod_{m=1}^{k}\frac{1}{w_{m}}\right)^{1/k}$. Given $\prod_{m=1}^{k}w_{m}=\frac{\prod_{m=1}^{k}l_{m}}{d_{i}}$, we obtain $\sum_{m=1}^{k}\frac{1}{w_{m}}\geq k\cdot\left(\frac{d_{i}}{\prod_{m=1}^{k}l_{m}}\right)^{1/k}$. Equality is achieved when all terms $\frac{1}{w_{m}}$ are equal, which implies $w_{1}=w_{2}=\cdots=w_{k}$. In practice, however, exact equality may not be attainable because each $\frac{l_{m}}{w_{m}}$ must be a natural number.
In the 2D task space example shown in Figure 8, the mapping for the $(18,12)$ space is optimal because the workload vector is $(w_{1},w_{2})=\left(\frac{l_{1}}{d_{1}},\frac{l_{2}}{d_{2}}\right)=(6,6)$, where the two components are equal. Similarly, the 3D example in Figure 9 uses a task space of $(4,8,4)$ mapped onto 16 processors. The solution shown in the figure yields a workload vector $(w_{1},w_{2},w_{3})=(2,2,2)$, which satisfies the sufficient condition for attaining the minimum communication volume.
These examples suggest that the optimal solution is achieved when the task space is divided among the processor dimensions in a balanced manner. However, in practice, this ideal may not be attainable due to the integrality constraint that each $\frac{l_{m}}{w_{m}}$ must be a natural number. We introduce a search-based algorithm for solving the optimization problem in Section 4.3.
As mentioned above, our analysis assumes locality in communication (e.g., such as nearest-neighbor communication), which is very common, but not universal, in practice. Other typical patterns, including anisotropic halo and all-to-all exchanges, are discussed in Section A.3 (see supplementary material). For algorithms without spatial locality or highly unbalanced tasks, decompose may be ineffective. In such cases, a good Mapple mapper may rely on other features, such as load-balancing (see Section 7).
4.3. Algorithm Implementation and Complexity Analysis
At a high level, the integrality constraint requires us to enumerate all possible ways of factoring $d_{i}$ into $k$ positive integers $d_{i_{1}},\ldots,d_{i_{k}}$ such that $\prod_{m=1}^{k}d_{i_{m}}=d_{i}$, and then select the factorization that minimizes the objective function $\sum_{m=1}^{k}\frac{d_{i_{m}}}{l_{m}}$. The main challenge lies in efficiently and exhaustively enumerating all valid factorizations to ensure optimality.
We begin with a simple case: suppose $k=3$ and $d_{i}=16=2^{4}$. To enumerate all possible factorizations, we must determine all ways to distribute the four factors of 2 across the three dimensions. This reduces to finding all non-negative integer solutions to the equation $x_{1}+x_{2}+x_{3}=4$, which can be solved using recursion or backtracking.
A more complex example is when $k=3$ and $d_{i}=48=2^{4}\cdot 3^{1}$. In this case, we must find all non-negative integer solutions to both $x_{1}+x_{2}+x_{3}=4$ and $y_{1}+y_{2}+y_{3}=1$, corresponding to the distribution of the prime factors 2 and 3, respectively. Each valid factorization corresponds to a tuple of the form $(2^{x_{1}}\cdot 3^{y_{1}},2^{x_{2}}\cdot 3^{y_{2}},2^{x_{3}}\cdot 3^{y_{3}})$ for some choice of $(x_{n},y_{n})$.
In the general case where $d_{i}=p_{1}^{a_{1}}\cdot p_{2}^{a_{2}}\cdot\ldots\cdot p_{t}^{a_{t}}$, the prime factorization naturally decomposes the enumeration task. We can independently enumerate the placement strategies for each prime factor by solving $t$ separate integer partition problems of the form $z_{1}+\ldots+z_{k}=a_{j}$ for each $p_{j}$, and then compute the Cartesian product of their solutions to generate all valid factorizations.
We now analyze the complexity of the search-based algorithm. Since the algorithm exhaustively enumerates all valid factorization strategies, its complexity is determined by the size of the search space. As shown in the earlier example, enumerating non-negative integer solutions to the equation $x_{1}+x_{2}+x_{3}=4$ is equivalent to counting integer partitions with repetition, which can be mapped to the problem of selecting two dividers among six gaps. This corresponds to the number of combinations $\binom{6}{2}=15$. In general, for a processor count with prime factorization $d_{i}=p_{1}^{a_{1}}\cdot p_{2}^{a_{2}}\cdots p_{t}^{a_{t}}$, the total number of factorization strategies is $\prod_{j=1}^{t}\binom{a_{j}+k-1}{k-1}$. Each term in the product counts the number of ways to distribute the $a_{j}$ copies of prime $p_{j}$ across $k$ dimensions. In practice, the exponents $a_{j}$ are usually small (typically less than 10), and the number of dimensions $k$ is often no greater than 3. As a result, the search space remains small, and the algorithm is computationally efficient in real-world scenarios.
We believe that enumeration is necessary to guarantee an optimal solution. Consider a greedy algorithm that, at each step, assigns prime factors of $d_{i}=72$ (i.e., $2,2,2,3,3$) to dimensions in a way that minimizes the maximum difference between elements of the workload vector. Given a task space of $(l_{1},l_{2})=(8,9)$, this greedy strategy produces a suboptimal workload vector of $\left(\frac{4}{3},\frac{3}{4}\right)$, which results in imbalanced partitioning. In contrast, our search-based algorithm identifies the optimal factorization that yields a perfectly balanced workload vector of $(1,1)$.
5. Implementation
We describe the translation of Mapple to the mapping interface of task-based runtime systems. These systems employ a highly asynchronous and pipelined execution model, where tasks progress through multiple stages to maximize throughput. Mapping decisions, which determine where tasks and data are placed, must be made at various points along this pipeline and are exposed via many distinct callback functions, each tied to a specific stage in a task’s lifecycle. This results in a fragmented and low-level interface that mirrors the internal structure of the runtime rather than providing an intuitive programming model for developers.
The core challenge is to design a high-level, unified mapping abstraction that can be faithfully translated into this low-level execution model. To illustrate this semantic gap, we first describe the runtime’s execution semantics in Section 5.1, then present our translation strategy in Section 5.2.
5.1. Execution Semantics of the Runtime System
Task-based runtime systems maximize throughput through an asynchronous, pipelined execution model, where a task advances through several distinct stages during its lifetime. To illustrate the low-level nature of the mapping interface, we present a much simplified version of the execution semantics of the Legion runtime system (bauer2012legion), focusing on the portion of the pipeline that determines the processor on which a task runs. Our semantics makes explicit a key source of complexity: the interface is tightly coupled to internal execution stages, where what is conceptually a single decision—choosing a processor for a task—is split across multiple pipeline stages at a low level of abstraction. The same low-level complexity exists in frameworks like StarPU (augonnet2010starpu), Charm++ (kale1993charm++), and HPX (kaiser2014hpx), which expose mapping through asynchronous callbacks and fragmented APIs. In contrast, Mapple provides a high-level DSL that centralizes this logic while still providing the same flexibility.
Figure 11 defines the structures used in the execution semantics. A task space is a set of multidimensional points, each representing a single point task to be executed. Working with sets of tasks is important for the efficiency of the runtime system itself, especially in the early stages of the task pipeline, where runtime operations can be done in constant time on a set rather than in linear time on each individual point task in the set.
The execution context consists of a set of tasks TT [^1] One major simplification is that we ignore dependencies between tasks; dependencies are not needed to illustrate the pipeline for choosing a task’s processor., a log of actions on tasks, and two vectors of $n$ queues, one queue per node of the machine. [^2] We also do not model nodes, using only processors, but in a full description a set of processors is associated with each node. When a task space is first enqueued in the runtime system, it goes into a distribution queue, from where its set of points may be subdivided and sent to other nodes for further processing. Eventually (perhaps after multiple recursive subdivisions of the set of tasks), a task space is moved to a mapping queue on a particular processor, where the assignment of the task space to a (potentially different) processor for execution will be made. This structure, where task spaces are subdivided and spread out to multiple processors for mapping, parallelizes the mapping process, preventing a single processor from becoming a runtime bottleneck. Once a task space is mapped, its point tasks are enqueued for execution on its assigned processor; because our focus is on the mapping process, we do not illustrate task execution here.
[…]
Figure 10. Abstract Syntax for Execution Model
$\frac{\begin{array}[]{l}t\in T\\ \text{enqueued}(t)\notin L\\ \end{array}}{L,E,M\xrightarrow{T}L\cup\{\text{enqueued}(t)\},E+_{1}t,M}$ [Enqueue]
$\frac{\begin{array}[]{l}\text{SHARD}(t)=\text{distribute}((t_{1},p_{1}),(t_{2},p_{2}))\\ t=(id,P),\ t_{1}=(id_{1},P_{1}),\ t_{2}=(id_{2},P_{2})\\ id_{1},id_{2}\text{ fresh (not in $T,E,M,L$)}\\ P=P_{1}\cup P_{2}\\ \end{array}}{L,E,M\xrightarrow{T}L,((E-_{p}t)+_{p_{1}}t_{1})+_{p_{2}}t_{2},M}$ [Distribute]
$\frac{\begin{array}[]{l}\text{SHARD}(t)=\text{local}(t)\\ t=(id,P)\\ E^{\prime}=E-_{i}t\end{array}}{L,E,M\xrightarrow{T}L,E^{\prime},M+_{i}t}$ [Local]
$\frac{\begin{array}[]{l}\text{SLICE}(t)=p\\ M^{\prime}=M-_{i}t\end{array}}{L,E,M\xrightarrow{T}L\cup\{\text{mapped}(t,p)\},E,M^{\prime}}$ [Map]
Figure 11. Formalization of the mapping pipeline.
Figure 11 formalizes the task execution rules described above using an operational semantics. Each judgment has the form
$L,E,M\xrightarrow{T}L^{\prime},E^{\prime},M^{\prime}$
which should be read as: given a set of task spaces $T$, the current execution log $L$ and vectors of queues $E,M$, transition to a new execution log $L^{\prime}$ and new vectors of queues $E^{\prime},M^{\prime}$. The rules rely on two user-supplied mapping functions, SHARD and SLICE, with the following signatures:
$\begin{array}[]{rl}\mbox{SHARD: }&t\rightarrow\mbox{distribute}((t\ast p)\ast(t\ast p))+\mbox{local}(t)\\ \mbox{SLICE: }&t\rightarrow p\end{array}$
The SHARD function takes a task space and either partitions it into smaller task spaces for distribution across nodes or assigns it to the local node. SLICE then selects a specific processor within the target node for execution. Although these are two separate callback functions in the low-level interface, Mapple unifies them by interpreting both as part of a single index transformation from task space to processor space.
Figure 11 formalizes the rules described above as execution judgments. The rules use an operator $_{i}$, which adds a task space to the $i$th queue of a vector of queues:
$(Q_{1},\ldots,Q_{i},\ldots Q_{n})+_{i}t=(Q_{1},\ldots,t:Q_{i},\ldots Q_{n})$
We also need the dual dequeuing operation:
$(Q_{1},\ldots,Q_{i}:t,\ldots Q_{n})-_{i}t=(Q_{1},\ldots,Q_{i},\ldots Q_{n})$
Note that dequeue is only defined if the task space $t$ is at the head of the selected queue.
We briefly summarize each rule. [Enqueue] submits a task space to the runtime system on node 1.333A simplification; in actual implementations, tasks can begin on any node.[Distribute] partitions a task space into smaller task spaces for mapping. [Local] selects the node where a task space’s mapping decision will be made using SHARD. [Map] assigns a task space to a specific processor within that node using the SLICE function.
The semantics of task execution underscore the low-level nature of mapping in task-based systems. To achieve high performance, task spaces progress through multiple pipelined stages, each requiring user-defined mapping decisions. For brevity, we focus on SHARD and SLICE, though in practice, 19 distinct callback APIs are invoked across a task’s lifetime. This design fragments mapping logic across numerous low-level callbacks, making the APIs complex and opaque to most application developers.
5.2. Translation
The translation of Mapple code to the low-level runtime interface poses two key challenges. First, as discussed in Section 5.1, Legion’s mapping logic is spread across many callback functions triggered at different stages of a task’s lifetime. While this structure mirrors the runtime’s internal pipeline, it forces developers to work with a scattered and unintuitive interface (Figure 1). Second, the low-level API lacks abstraction over the machine architecture, i.e., there is no logical view of the hardware, nor support for aligning mapping decisions with task spaces.
We implement Mapple by targeting Legion’s programmatic mapping interface, comprising roughly 7,000 lines of C++ code—much of it dedicated to handling low-level runtime details. Mapple overrides a set of callback functions invoked at different stages of the task execution pipeline to decide task placement. The two core overrides, corresponding to the user-defined SHARD and SLICE functions, are illustrated in Figure 1 and explained in Section 5.1.
We now give a concrete example using the block1D_x distribution (Figure 7). Let the processor space m = Machine(GPU) have shape $(n,g)$, representing $n$ nodes and $g$ GPUs per node. The mapper applies $m_{1}=m.\mathtt{merge}(0,1).\mathtt{split}(0,ng)$, which reshapes the processor space to $(ng,1)$. For an index point $(i_{x},i_{y})$ in a task space of size $(L_{x},L_{y})$, the block1D_x function returns point $(\lfloor i_{x}\times\tfrac{ng}{L_{x}}\rfloor,0)$ in $m_{1}$. The Mapple implementation then converts this point in $m_{1}$ back to a point in the original 2D processor space $m$ by applying the processor space transformations in reverse order (i.e., first inverting $\mathtt{split}(0,ng)$ and then inverting $\mathtt{merge}(0,1)$). Thus, the $m_{1}$ index $(\lfloor i_{x}\times\tfrac{ng}{L_{x}}\rfloor,0)$ maps back to a point in $m$ as $(\text{node\_id},\text{gpu\_id})=(a\bmod n,\lfloor a/n\rfloor)$ where $a=\lfloor i_{x}\times\tfrac{ng}{L_{x}}\rfloor$. In the original space of $m$, node_id identifies the node and gpu_id specifies the GPU within that node. The node index is used by SHARD, applying either [Distribute] or [Local] depending on whether the target node is local, while the GPU index is used by SLICE under the [Map] rule.
6. Results
Experiments were run on a cluster where each node has 40 IBM Power9 CPU cores and 4 NVIDIA V100 GPUs connected with NVLink 2.0 and an Infiniband EDR interconnect. To account for performance variability, each measurement was repeated five times, and the average is reported.
Our approach is evaluated on a suite of nine benchmarks chosen to demonstrate the generality and effectiveness of Mapple across representative distributed workloads. We use the published artifact repositories (yadav2022distal; bauer2012legion) and their corresponding problem sizes, where the mappers were developed by experts and tuned for the same machine, representing state-of-the-art implementations.
The first three benchmarks are drawn from scientific simulation workloads. Circuit (bauer2012legion) simulates electrical circuit behavior by modeling currents and voltages across interconnected nodes and wires. Stencil (stencilvan2014parallel) operates on a 2D grid, updating each point’s value based on a stencil pattern derived from its neighbors. Pennant (ferenbaugh2015pennant) models unstructured mesh Lagrangian staggered-grid hydrodynamics, essential for simulating compressible flow.
The remaining six benchmarks are parallel matrix multiplication algorithms, an extensively studied problem in parallel computing that underpins many scientific applications. Matrix multiplication also serves as a classic stress test for distributed systems, as it requires carefully balancing computation and communication across multiple processors. They matrix multiplication algorithms we consider are Cannon’s (cannon1969cellular), SUMMA (van1997summa), PUMMA (choi1994pumma), Johnson’s (johnsonagarwal1995three), Solomonik’s (solomonik2011communication), and COSMA (cosmakwasniewski2019red). They are divided into 2D and non-2D categories. The 2D algorithms, such as Cannon’s, PUMMA, and SUMMA, enhance communication efficiency by partitioning matrices into 2D tiles and mapping them onto the processor space, utilizing techniques like pipelining. Conversely, non-2D algorithms like Johnson’s, Solomonik’s, and COSMA, employ 3D partitioning or optimize processor grids to minimize communication overhead and boost performance.
6.1. Lines of Code Reduction
[…]
We manually verify that both approaches produce identical mapping decisions and matching performance (i.e., identical throughput), showing that any overhead introduced by Mapple is negligible. This confirms the correctness and efficiency of our Mapple mappers relative to the original low-level C++ implementations in the published artifact repositories (yadav2022distal; bauer2012legion).
We note that the application code is substantially larger than the mapping code and remains unchanged. The three scientific computing applications, Circuit, Stencil, and Pennant, are implemented in Regent (slaughter2015regent) with 776, 825, and 3.6K lines, respectively, and the six distributed matrix multiplication benchmarks each contain about 7.5K lines of Legion C++ code.
[…]
Application 1 2 3 4 5 6 7 8 9 Avg.
LoC in C++ 347 306 379 447 437 430 428 433 448 406
LoC in Mapple 16 14 16 38 38 38 33 38 32 29
[…]
6.2. Performance Tuning
[…]
This performance gain comes from Mapple’s high-level declarative design, which is still expressive for high-performance mapping. Although our compiler translates Mapple mappers into standard C++ code, making the same strategies theoretically possible in C++, experts have not discovered them in practice. This is likely due to the low-level complexity of C++ mappers, where exploring new strategies often requires rewriting hundreds of lines of code. In contrast, similar changes in Mapple take only a few lines, making tuning much easier.
We further analyze the sources of performance gains. For the six matrix multiplication applications (indexed 4 to 9), improvements arise solely from optimized index mappings. For the three scientific applications, they result from different memory assignments enabled by Mapple’s full feature set (Section 7). Finally, case studies in Section A.5 (see supplementary material) illustrate how mapping choices impact performance.
Table 2. Performance improvements of Mapple mappers over expert-designed C++ mappers.
Application 1 2 3 4 5 6 7 8 9
[…]
Figure 12. The distribution of improvement percentage over different configurations.
Figure 13. Geometric mean of improvement percentage w.r.t. aspect ratios of task spaces.
Figure 14. Geometric mean of improvement percentage w.r.t. area of task space per node.
Figure 15. Geometric mean of improvement percentage w.r.t. machine sizes.
6.3. Performance Study of the Decompose Primitive
To evaluate the impact of the decompose primitive, we compare it with the default machine space reshaping heuristic from Algorithm 1 (see supplementary material). We measure end-to-end performance on stencil applications, which are common in image processing and scientific computing, where each point updates based on its neighbors. As these applications are sensitive to data partitioning and mapping, the results provide strong evidence of decompose improving communication efficiency in practical workloads.
To ensure a fair comparison, we systematically vary the 2D stencil parameters. We test six aspect ratios ($1:1$, $1:2$, $1:4$, $1:8$, $1:16$, $1:32$), representative of shapes typical in scientific simulations (e.g., liquid films (nave2010direct), shock tubes (wong2019high)). We also vary the task space size per node across five scales ($10^{6}$, $10^{7}$, $10^{8}$, $2{\times}10^{8}$, $4{\times}10^{8}$) for different communication-to-computation ratios, and scale the number of GPUs from 4 to 128 in powers of two. In total, we evaluate 180 configurations.
We report percentage improvements in Figure 15, ranging from 0% to 83%, with a geometric mean of 16%. Figure 15 plots the geometric mean improvement against the task space aspect ratio. As the ratio increases from $1:1$ to $1:32$, the improvement rises from 7% to 27%, showing the growing benefit of using the decompose primitive. This trend is expected since Algorithm 1 evenly splits processors, which works best when the task space is nearly square.
Figure 15 shows how improvement varies with task space size per node. As the area grows from $10^{6}$ to $4\times 10^{8}$, the improvement decreases from 32% to 5%, since larger spaces lower the communication-to-computation ratio. Figure 15 plots improvement against machine size. The benefit peaks at 26% on 4 nodes: increasing from 1 to 4 nodes amplifies inter-node communication, while beyond 4 nodes, inter-rack latency becomes the bottleneck, reducing the gain.
7. Discussion
Program $::=\;\text{Statement}^{+}$
Statement $::=\;\text{TaskMap}\mid\text{DataMap}\mid\text{DataLayout}\mid{}$
$\hskip 17.00024pt\text{FuncDef}\mid\mathrm{IndexTaskMap}\;\mathit{t}\;\mathit{v}\mid\cdots$
TaskMap $::=\;\mathrm{Task}\;\mathit{t}\;\text{Proc}^{+}$
DataMap $::=\;\mathrm{Region}\;\mathit{t}\;\mathit{r}\;\text{Proc}\;\text{Memory}^{+}$
Proc $::=\;\mathrm{CPU}\mid\mathrm{GPU}\mid\mathrm{OMP}$
Memory $::=\;\mathrm{SYSMEM}\mid\mathrm{FBMEM}\mid\mathrm{ZCMEM}$
DataLayout $::=\;\mathrm{Layout}\;\mathit{t}\;\mathit{r}\;\text{Proc}\;\text{Constraint}^{+}$
Constraint $::=\;\mathrm{SOA}\mid\mathrm{AOS}\mid\mathrm{C\_order}\mid\mathrm{F\_order}\mid\mathrm{Align}{\ ==\ }\mathit{i}$
FuncDef $::=\;\mathrm{def}\;\mathit{v}(\mathit{v}^{+}):\;\text{FuncStmt}^{+}$
FuncStmt $::=\;\mathit{v}\;{=}\;\text{Expr}\mid\mathrm{return}\;\text{Expr}$
Expr $::=\;\mathit{v}\mid\mathit{v}(\text{Expr}^{+})\mid\mathrm{Machine}(\text{Proc})\mid{}$
$\hskip 17.00024pt\text{Expr}\mathbin{.}\text{Expr}\mid\text{Expr}\;\mathrm{Op}\;\text{Expr}\mid(\text{Expr})\mid{}$
[…]
Primitive
Primitive $::=\;\mathrm{split}\mid\mathrm{merge}\mid\mathrm{swap}\mid\mathrm{slice}\mid\mathrm{decompose}$
$\mathit{t}\in\text{{Task Names}}\quad\mathit{r}\in\text{{Region Names}}\quad\mathit{v}\in\text{{Variables}}\quad\mathit{i}\in\text{{Integers}}$
Figure 16. Syntax of the Mapple language.
While Section 3 and Section 4 focus on index mapping, the full optimization space includes several additional dimensions that contribute to the performance gains observed in the three scientific applications evaluated in Section 6.2. Figure 16 presents a simplified grammar capturing the core constructs of Mapple. A primary axis of control is processor selection, expressed via the TaskMap directive. This determines whether a given task is assigned to GPUs, CPUs, or the OpenMP runtime. The decision is made per task and depends on factors such as task granularity, GPU memory capacity, and kernel launch overhead. For example, small tasks may favor CPUs to avoid the cost of GPU launches, and large-memory tasks may also favor CPUs if they exceed GPU memory limits.
Another critical dimension is memory placement, specified through the DataMap construct. This governs where task arguments are stored: in GPU FrameBuffer memory for high-speed access, in ZeroCopy regions to enable CPU-GPU sharing, or in host memory when capacity is a concern. Each location involves trade-offs between access latency, available memory, and transfer costs. This mapping is defined per task and per argument.
The optimization space also includes memory layout, configured with the DataLayout statement. This involves selecting between layouts such as Struct of Arrays (SoA) versus Array of Structures (AoS), specifying memory ordering (e.g., Fortran vs. C order), and applying alignment constraints (e.g., 128-byte alignment). These layout choices directly impact cache behavior and performance. This decision is made per task, per data region, and per target processor. Additional features, not shown in the figure, include scheduling policies (e.g., task prioritization), garbage collection strategies, and load-balancing directives.
While Mapple captures a broad range of optimization dimensions, it is intentionally constrained to focus on performance-critical mapping decisions that, in our experience, are the most common in practice. One current limitation is that Mapple is stateless, making it less suitable for highly dynamic workloads or computations that would benefit from retaining historical information across mapping decisions. Although it could be extended with a persistent state, we have not yet found this necessary. Beyond this, we have not encountered cases where Mapple is less expressive than the C++ interface. Notably, every mapper that achieved higher performance than the expert-written C++ implementations was expressible in Mapple, demonstrating that Mapple is effective despite being more restrictive than general-purpose C++, at least for the applications we studied.
8. Related Work
8.1. Task-based Programming Systems
Table 3. Mapping features exposed to application control by different systems and covered by Mapple.
Systems Task Placement Data Placement Data Layout Scheduling Load Balancing
Legion (bauer2012legion)
StarPU (augonnet2010starpu)
Chapel (chamberlain2007parallelchapel)
Charm++ (kale1993charm++)
PaRSEC (danalis2015parsec)
X10 (charles2005x10)
HPX (heller2017hpx)
OpenMP (chandra2001parallelopenmp)
OmpSs (duran2011ompss)
Mapple
Table 3 summarizes mapping features in HPC systems, all of which are supported by Mapple. Systems like Legion (bauer2012legion) and StarPU (augonnet2009starpu) offer low-level C/C++ mapping interfaces via asynchronous callbacks, requiring deep familiarity with internal runtime abstractions. Other HPC systems offer less direct mapping control. ZPL (deitz2004abstractionszpl), Chapel (chamberlain2007parallelchapel), and X10 (charles2005x10) expose predefined data distributions (e.g., blocked, cyclic). HPX (heller2017hpx) and PaRSEC (danalis2015parsec) rely on schedulers with tunable policies for heterogeneous execution. OpenMP (chandra2001parallelopenmp), OmpSs (duran2011ompss), Charm++ (kale1993charm++), and Chapel’s low-level APIs allow programmers to specify task placement directly in the source code, ranging from processor annotations (OpenMP, OmpSs) to custom C++ mapping classes (Charm++). Task-based systems in data analytics and AI such as MapReduce (dean2008mapreduce), Spark (zaharia2010spark), Dask (rocklin2015dask), Ray (moritz2018ray), and Pathways (barham2022pathways) rely on runtime mapping heuristics. This eases development but limits performance tuning when heuristics fall short. Mapple offers better ease-of-use than low-level mapping interfaces while enabling finer-grained control than systems that rely entirely on heuristics built into the runtime system.
8.2. Loop Transformations
A large body of work on compiler algorithms uses loop transformations to maximize parallelism and minimize communication, including unimodular transformations (banerjee1990unimodular; banerjee2007loop; wolf1992improving; wolf1991loop; wolfe1982optimizing) (i.e., combinations of loop interchange, reversal, and skewing), loop block/tiling, fusion, fission, reindexing, and scaling (lim1999affine; lim2001blocking; wolf1991data; xue2000loop), and affine transformations used by the polyhedral model (bondhugula2013compiling; bondhugula2008practical; acharya2018polyhedral). These works share a similar motivation to Mapple, though the goals and specific transformations are different as explained in Section 1.
Another line of work uses scheduling languages (chen2008framework; ragan2012decoupling; zhang2018graphit; baghdadi2019tiramisu; kjolstad2017tensor; senanayake2020sparse; chen2018tvm; yadav2022spdistal) to separate the algorithm from the schedule, based on traditional loop optimizations such as split, collapse, and loop reordering. This separation enables programmers to just change the schedule when moving to different hardware. A growing body of research focuses on auto-scheduling (adams2019learning; chen2018learning; mullapudi2016automatically; zheng2020ansor; zheng2022amos; tollenaere2023autotuning) to reduce the manual effort required in writing schedules. Task-based systems share a similar philosophy of separating the algorithm from its schedule or mapping. However, prior scheduling languages such as DISTAL (yadav2022spdistal) and Distributed Halide (denniston2016distributed) are specialized compiler frameworks for tensor algebra that lower high-level tensor expressions to distributed execution. In contrast, Mapple addresses the general mapping problem in task-based runtime systems (see Section 7), supporting a broader class of task-based programs. Its design centers on the decompose primitive (see Section 4), which automatically resolves dimensionality mismatches between task and processor spaces. In our evaluation, DISTAL’s C++ mappers serve as baselines; Mapple reduces their code size by over 11× and discovers mappers that achieve higher performance.
9. Conclusion
[…]
Appendix A Supplementary Material
A.1. Example of Cannon’s Algorithm
Cannon’s algorithm computes $A=B\cdot C$ by arranging all processors in a 2D grid, and assigning a tile of each matrix to each of the processors. Every processor $p$ alternates between systolic shifts of the tiles of $B$ and $C$ with $p$’s row and column neighbors and accumulating the product of $p$’s current $B$ and $C$ tiles into $A$.
We illustrate the task-based pseudocode for Cannon’s algorithm, adapted from code generated by DISTAL (yadav2022distal). The program in Figure 1(a) partitions the input matrices into tiles and launches the tiles task across the processor grid. Each tiles task, in turn, launches multiple systolic tasks that fetch tiles of $B$ and $C$ and accumulate their products into the corresponding tile of $A$. This program builds a task graph for distributed matrix multiplication, where data dependencies between tasks are automatically managed by the runtime system. The task space of the tiles task is two-dimensional, defined as $(i,j)\in\{0,\ldots,p_{x}-1\}\times\{0,\ldots,p_{y}-1\}$, where each point corresponds to a tile of the output matrix. The mapping from this task space to the processor space is specified by the mapper declared in the IndexTaskMap statement shown in Figure 1(b).
1task systolic(A: region(double),
2 B: region(double),
3 C: region(double))
4 writes(A), reads(A, B, C):
5 # Internally use cuBLAS for single-GPU matmul.
6 A += matmul(B, C)
7
8task tiles(A: region(double),
9 B: region(double),
10 C: region(double),
11 pB: partition(B),
12 pC: partition(C),
13 px: int, i: int, j: int)
14 writes(A), reads(A, B, C):
15 for k in (0, px):
16 k = (k + i + j) % px
17 systolic(A, pB[i, k’], pC[k’, j])
18
19task cannons_matmul(A: region(double),
20 B: region(double),
21 C: region(double),
22 px : int, py : int)
23 writes(A), reads(A, B, C):
24 pA, pB, pC = ... # Partition into px * py tiles.
25 for i, j in ({0, 0}, {px, py}):
26 tiles(pA[i, j], B, C, pB, pC, px, i, j)
(a) Application: pseudocode for Cannon’s algorithm.
1# Define a block mapping for a 2D launch domain.
2def block(Task task):
3 # ispace represents 2D launch domain of a x b.
4 is = task.ispace
5 # ipoint is an index point within launch domain.
6 ip = task.ipoint
7 # Get all GPUs as m nodes x n GPUs grid.
8 machine = Machine(GPU)
9 # Tile nodes; 3-dim grid: (m1 x m2) x n.
10 m3d = machine.decompose(0, is)
11 # Tile GPUs; 4-dim grid: (m1 x m2) x (n1 x n2).
12 m4d = m3d.decompose(2, is / m3d[:-1])
13 # Project the index point onto the 4D machine.
14 upper = (ip[0] * m4d[0] / is[0],
15 ip[1] * m4d[1] / is[1])
16 lower = (ip[0] % m4d[2], ip[1] % m4d[3])
17 # Unpack two 2D tuples to index 4D machine.
18 return m4d[*upper, *lower]
19
20IndexTaskMap tiles block
21
22# Map all regions of all tasks to GPU framebuffer.
23Region * * GPU FBMEM
24# Data layout in memory.
25Layout * * * C_order
26# Collect read-only regions used by task_4.
27GarbageCollect systolic *
28# At most 1 systolic task is active at a time.
29Backpressure systolic 1
(b) Mapper in Mapple for Cannon’s algorithm.
Figure A1. Application code in a task-based programming system and the corresponding Mapple mapper for the Cannon’s algorithm.
The mapper that faithfully implements Cannon’s algorithm organizes the processors into a 2D grid and blocks the GPUs on each node. Here, we use the decompose machine transformation twice to block the node and GPU dimensions of the machine to resolve the dimensionality mismatch between the task space and the processor space. Lines 2329 specify the mapping of data to memories. The Regent statement specifies that all data should be mapped to GPU framebuffer memory. The Layout directive declares that any region used by any task on any processor (denoted by three *s) is allocated in row-major (C-style) order, which is necessary to interface with cuBLAS. The GarbageCollect directive ensures the read-only tiles of $B$ and $C$ requested by systolic tasks are eagerly reclaimed to minimize memory usage. Finally, the Backpressure directive limits the number of active systolic tasks to 1 per processor to avoid allocating too many tiles at one time.
A.2. Baseline Greedy Algorithm
1
Input:
   $d$ # Number of processors
   $k$ # Dimensionality of the task space
2
3Function Greedy($d$, $k$):
4  primes $\leftarrow$ PrimeFactorization ($d$) # Sorted list of prime factors, $d=p_{1}\leq\dots\leq p_{n}$
5  factors $\leftarrow$ new int[$k$]
6  for $i\in\{0,\dots,k{-}1\}$ do
7     factors[$i$] $\leftarrow 1$
8 for $p\in$ primes do
9     $j\leftarrow$ ArgMin (factors) # Index of smallest current product
10     factors[$j$] $\leftarrow$ factors[$j$] $\times\ p$
11 Sort (factors, descending=True) # Sort for consistent ordering
12  return factors
Algorithm 1 Greedy Heuristic for Processor Grid Selection (Suboptimal)
A.3. Generalization of the Decompose Primitive
While Section 4 addresses isotropic communication across task space dimensions, the decompose primitive can be extended to anisotropic patterns, such as uneven halo widths and dimension-specific all-to-all exchanges (e.g., for data transposes). This generalization supports a much wider range of computations on block-structured grids. In both cases, only the objective function changes; the same minimization procedure from Section 4.3 still applies.
A.3.1. Anisotropic halo communication
To generalize, we extend the communication area from Section 4.2 to a volume that captures the total data exchanged between distributed machines. In isotropic cases, where halo width is uniform, volume reduces to area. For anisotropic patterns, varying halo width across dimensions must be considered. Letting $h_{n}$ denote the width in dimension $n$, the total communication volume $V$ in a $k$-dimensional space is $V=\sum_{n=1}^{k}d_{i_{n}}h_{n}\prod_{m\neq n}l_{m}$. For $l_{n}=w_{n}d_{i_{n}}$, the volume, $V$, can also be expressed as $V=\left(\sum_{n=1}^{k}\frac{h_{n}}{w_{n}}\right)\left(\prod_{m=1}^{k}l_{m}\right)$, where $\prod_{m=1}^{k}l_{m}$ is constant for a certain task space.
A.3.2. Transpose via all-to-all communication along certain dimensions
For transpose operations via all-to-all communication, the data volume per partition along the $n$-th dimension is given by $v_{n}=\frac{d_{i_{n}}-1}{d_{i_{n}}}\prod_{m=1}^{k}w_{m}$. This reflects that each partition splits its data evenly into $d_{i_{n}}$ groups, one of which remains local while the others are sent to distinct peers along the $n$-th dimension. Thus, the total communication volume within a pencil of $d_{i_{n}}$ aligned processors is $d_{i_{n}}v_{n}$. To optimize the mapping, we minimize the total communication volume, $\left(V+\sum_{n\in\mathbb{T}}V_{n}^{*}\right)$, where $\mathbb{T}$ denotes the set of dimensions requiring transposes. The volume along dimension $n$, $V_{n}^{*}$, is defined as:
$V_{n}^{*}=v_{n}\prod_{m=1}^{k}d_{i_{m}}=\left(1-\frac{1}{d_{i_{n}}}\right)\left(\prod_{m=1}^{k}w_{m}\right)d_{i}.$
A.4. Example Mapping Functions in Mapple
Figure A2. Example functions used by the mappers for the distributed matrix-multiplication algorithms.
A.5. Case Studies on Mapper Performance
Here we present a case study highlighting the critical role of precise control over mapping strategies. Our investigation demonstrates significant performance variations resulting from minor alterations in the hierarchical_block2D function for the Cannon’s, PUMMA, and SUMMA algorithms. Specifically, we contrast the mapping function specified by the algorithm with an alternative approach integrating runtime heuristics: the runtime system dynamically assigns the index point to one of the four GPUs on the node, selecting the GPU with the least workload at runtime, rather than adhering to a predetermined distribution.
[…]
(a) Cannon’s
(b) PUMMA
(c) SUMMA
Figure A3. Throughput comparison between the original hierarchical_block2D mapping function (specified by the Cannon’s, PUMMA, and SUMMA algorithm) and a runtime heuristics-based mapper. The runtime heuristics-based mapping can lead to out-of-memory error (denoted by “OOM”).